OpenStudy (anonymous):

33 less than twice a number is equal to the number. What is the number?

OpenStudy (amistre64):

2n - 33 = n

OpenStudy (anonymous):

\[2x - 33 = y\]

OpenStudy (anonymous):

2x-33 = x 2x-x=33 x =33

OpenStudy (anonymous):

\[2x = y+33\]

OpenStudy (amistre64):

rgupta; close, but then you have to take that a step further and say that x=y otherwise its solution is a line

OpenStudy (anonymous):

let number = x twice of the number = 2x 2x - 33 = x 2x-x = 33 x = 33

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

33 is right but i have seen some doozies here today. i guess caveat emptor